# Sharpness of Minima in Deep Matrix Factorization: Exact Expressions

### v1

@@ 1 Introduction
Related Work Depth Input Dim. Output Dim. $\lambda_{max}(\nabla^{2}\mathcal{L}(\bm{w}))$ Layers
(Mulayoff & Michaeli, 2020) $L$ $d_{0}$ $d_{L}$ $\bm{w}\in\Omega_{0}$ $\mathbb{R}^{a\times b}$
(Zhu et al., 2023) 22 11 11 $\bm{w}\in\mathbb{R}^{N}$ $\mathbb{R}$
(Singh & Hofmann, 2024) 22 11 11 $\bm{w}\in\mathbb{R}^{N}$ $\mathbb{R}^{a}$
(Ghosh et al., 2025) $L$ $d_{0}$ $d_{L}$ $\bm{w}\in\Omega_{B}$ $\mathbb{R}^{a\times b}$
@@ 4 Overparameterized Deep Matrix Factorization
$2\sigma_{\max}(\bm{I}\otimes\bm{L}\bm{L}^{\top}+\bm{R}\bm{R}^{\top}\otimes\bm{I})=2\sigma_{\max}(\bm{I}\otimes\bm{L}\bm{L}^{\top})+2\sigma_{\max}(\bm{R}\bm{R}^{\top}\otimes.\bm{I}).$ (42)
@@ B.1 Proof of Theorem 5
$\lambda_{\max}(\mathcal{L}(\bm{w}^{*}))$ $=\max_{\begin{subarray}{c}\bm{U}_{1},\bm{U}_{2},\dots,\bm{U}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{U}_{i}}_{F}^{2}=1\end{subarray}}2\norm{\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\bm{U}_{i}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)\Biggr]}_{F}^{2}$ (67)
@@
$\lambda_{\max}(\mathcal{L}(\bm{w}^{*}))$ $=\max_{\bm{u}:\norm{\bm{u}}_{2}=1}2\norm{\bm{K}\bm{u}}_{2}^{2}$ (73)

### v2

@@ 1 Introduction
Related Work Depth Input Dim. Output Dim. $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}))$ Layers
Mulayoff & Michaeli (2020, Theorem 1) $L$ $d_{0}$ $d_{L}$ $\bm{w}\in\Omega_{0}$ $\mathbb{R}^{a\times b}$
Zhu et al. (2023, Appendix B.1) 22 11 11 $\bm{w}\in\mathbb{R}^{N}$ $\mathbb{R}$
Singh & Hofmann (2024, Theorem 1) 22 11 11 $\bm{w}\in\mathbb{R}^{N}$ $\mathbb{R}^{a}$
Ghosh et al. (2025, Lemma 1) $L$ $d_{0}$ $d_{L}$ $\bm{w}\in\Omega_{B}$ $\mathbb{R}^{a\times b}$
@@ 4 Overparameterized Deep Matrix Factorization
$2\sigma_{\max}(\bm{I}\otimes\bm{L}\bm{L}^{\top}+\bm{R}\bm{R}^{\top}\otimes\bm{I})=2\sigma_{\max}(\bm{I}\otimes\bm{L}\bm{L}^{\top})+2\sigma_{\max}(\bm{R}\bm{R}^{\top}\otimes\bm{I}).$ (42)
@@ B.1 Proof of Theorem 5
$\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ $=\max_{\begin{subarray}{c}\bm{U}_{1},\bm{U}_{2},\dots,\bm{U}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{U}_{i}}_{F}^{2}=1\end{subarray}}2\norm{\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\bm{U}_{i}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)\Biggr]}_{F}^{2}$ (67)
@@
$\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ $=\max_{\bm{u}:\norm{\bm{u}}_{2}=1}2\norm{\bm{K}\bm{u}}_{2}^{2}$ (73)